# Implicit Neural Shape Optimization for 3D High-Contrast Electrical Impedance Tomography

### v1

@@ 3.2 Optimization in Latent Space
Following [38], we assume the regularity condition that $\nabla_{x}f_{\theta}(z,x)\neq 0$ whenever $f_{\theta}(z,x)=0$, ensuring a well-defined surface with consistent normal vectors. This condition is naturally satisfied by properly trained neural SDF models that accurately approximate true signed distance functions.
This latent representation fundamentally transforms the optimization landscape by dramatically reducing the dimensionality of the problem — from the infinite-dimensional space of all possible interfaces to a compact latent space, typically of dimension $d\approx 256\ll\infty$. Such a dimensionality reduction not only enhances computational efficiency but also provides implicit regularization that helps mitigate the severe ill-posedness inherent in EIT reconstruction. By constraining the search space to the manifold of plausible shapes encoded by the pre-trained model, we effectively incorporate prior knowledge about realistic shape variations without requiring explicit regularization terms.
Figure 3: Integration of NDF with EIT reconstruction: (a) Pre-trained NDF model encodes anatomical shapes into a low-dimensional manifold $\mathcal{M}$ via diffeomorphic flows. (b) Our physics-informed optimization in $\mathcal{M}$ minimizes the EIT data misfit while inheriting NDF’s topology preservation properties.
3.3 Latent Space Gradient for Shape Optimization
To solve the optimization problem (13) efficiently, we derive analytical expressions for the gradient of the objective function with respect to the latent variables. This requires extending the classical shape calculus to the neural implicit setting. Let $V:\mathbb{R}\times\mathbb{R}^{3}\mapsto\mathbb{R}^{3}$ be a velocity field and let $x(t,X)$ be the solution to
@@ 3.3 Latent Space Gradient for Shape Optimization
Let $\Omega_{z}$ be a bounded domain of class $C^{2}$, and let $\mathcal{L}(z)$ be defined in (13). If $\nabla_{x}f_{\theta}$ does not vanish on $\Gamma_{z}$, then:
@@
This elegant gradient formulation reveals the intricate relationship between latent space variations and interface geometry changes. The integrand in (16) combines three critical components: (1) the sensitivity of the forward solution $\frac{\partial u}{\partial n}$, (2) the sensitivity of the adjoint solution $\frac{\partial w}{\partial n}$, and (3) the shape-to-latent mapping $\frac{\nabla_{z}f_{\theta}}{\|\nabla_{x}f_{\theta}\|}$. This formulation provides a mathematically rigorous bridge between neural network parameters and geometric optimization, allowing efficient navigation of the latent manifold toward solutions that match the physical measurements.
Building on this established gradient expression, we can update z using a stochastic gradient descent algorithm. In our experiments, we adopt the Adam method, which is an efficient optimization method based on gradient and moment, although other gradient descent methods could also be employed. The complete algorithm description is presented in Algorithm 1. Our approach offers several key advantages, including efficient optimization in a low-dimensional latent space and topology preservation inherited from the pre-trained neural distance function (NDF). This combination enhances the overall performance and effectiveness of our method, particularly for challenging 3D high-contrast EIT reconstruction problems, where maintaining correct topology is crucial for accurate interpretation.
Algorithm 1 Implicit Neural Shape Optimization for High-Contrast EIT
1:Input: Voltage measurements $\{V_{\text{meas}}\}$, pre-trained NDF model $f_{\theta}$.
2:Initialize: Latent code $z_{0}$, learning rates $\{\alpha_{k}\}$, max iterations $N$.
3:for $k=0$ to $N$ do
@@ 4 Convergence Analysis
$d^{2}J(\Gamma;V_{1};V_{2})=\lim_{t\rightarrow 0^{+}}\frac{dJ(\Gamma_{t}(V_{1});V_{2}(t))-dJ(\Gamma;V_{2}(0))}{t}.$ (18)
Following [11, Theorem 6.2, Chapter 9], we have
@@
$d^{2}J(\Gamma;V_{1}(0);V_{2}(0))$ $:=\lim_{t\rightarrow 0^{+}}\frac{dJ(T_{t}(\Gamma,V_{1});V_{2}(0))-dJ(\Gamma;V_{2}(0))}{t},$ (20)
@@
For a detailed discussion of second-order shape derivatives, we refer to [11, 35] and [20, Appendix].
@@
Let $\Gamma$ be of class $C^{5}$. Then there exists $C>0$ independent of $V$ such that:
$|dJ(\Gamma;V(0))|$ $\leq C\|V(0)\|_{C(\Gamma;\mathbb{R}^{3})},\quad\forall V\in C([0,\varepsilon);C(\overline{\Omega},\mathbb{R}^{3}))$ (24)
$|d^{2}J(\Gamma;V_{1}(0);V_{2}(0))|$ $\leq C\|V_{1}(0)\|_{C^{3}(\Gamma,\mathbb{R}^{3})}\|V_{2}(0)\|_{C^{3}(\Gamma,\mathbb{R}^{3})},$ (25)
for all $V_{1},V_{2}\in C([0,\varepsilon);C^{4}(\overline{\Omega},\mathbb{R}^{3}))$, where $\varepsilon>0$ is sufficiently small.
@@
Let $\Omega\subset\mathbb{R}^{n}$ be a bounded Lipschitz domain with boundary $\partial\Omega=\Gamma\cup\Sigma$. Consider the following problem
@@
By the integral by parts, we can easily get
[…]
Then using Cauchy-Schwarz and trace theorems, we have
@@
The modified Poincaré inequality gives that
@@
which maps an admissible boundary $\Gamma$ to the value of state function $u$ at $x_{0}\in\Sigma$. By [1, Theorem 2], $F$ is twice continuously differentiable with
@@
with boundary data
@@
$+(V_{1,\tau}\tau\cdot\nabla V_{2,n}+V_{2,\tau}\tau\cdot\nabla V_{1,n})\partial_{n}u,$ (34)
where $u$ is the solution of the state problem, and $n$ denotes the unit outward normal to the boundary $\Gamma$. The notation $V_{j,n}=V_{j}(0)\cdot n$ and $V_{j,\tau}=V_{j}(0)\cdot\tau$ denote the normal and tangential components of the vector $V_{j}(0)$, respectively, and H denotes the mean curvature of the boundary $\Gamma$. Moreover, $u_{j}^{\prime}$ for $j=1,2$, is the solution to the following boundary value problem
$\Delta u^{\prime}&=0\quad\text{ in }\Omega,\\ u^{\prime}&=V_{n}\partial_{n}u\quad\text{ on }\Gamma,\\ \partial_{n}u^{\prime}&=0\quad\text{ on }\Sigma.\end{aligned}\right.$ (35)
Still by [1][Theorem 1], we have
$dF(\Gamma;V_{1}(0);V_{2}(0))=u^{\prime}(x_{0}).$ (36)
@@
$|dF(\Gamma;V(0))|$ $\leq C\|V(0)\|_{C(\Gamma;\mathbb{R}^{3})},\quad\forall V\in C([0,\varepsilon);C(\overline{\Omega},\mathbb{R}^{3})),$ (37)
$|d^{2}F(\Gamma;V_{1}(0);V_{2}(0))|$ $\leq C\big{\|}V_{1}(0)\|_{C^{3}(\Gamma,\mathbb{R}^{3})}\|V_{2}(0)\|_{C^{3}(\Gamma,\mathbb{R}^{3})},\,\quad\forall V_{1},V_{2}\in C([0,\varepsilon);C^{4}(\overline{\Omega},\mathbb{R}^{3})).$ (38)
@@
First, the $C^{5}$ regularity of the boundary $\Gamma$ implies that, for any bounded subdomain $\Omega^{+}=\mathbb{R}^{3}\backslash\overline{\Omega}$,
$u\in H^{5}(\Omega^{+}),\quad\gamma u\in H^{9/2}(\Gamma)\quad\mbox{and}\quad\partial_{n}u\in H^{7/2}(\Gamma).$ (39)
@@
$|dF(\Gamma;V(0))|=|u^{\prime}(x_{0})|\leq\max_{x_{\in}\Sigma}|u^{\prime}(x)|=\|u^{\prime}\|_{C^{0}(\Sigma)}.$ (40)
@@
$\|u^{\prime}\|_{C^{0}(\Sigma)}\leq C\|u^{\prime}\|_{H^{3/2}(\Sigma)}\leq C\|u^{\prime}\|_{H^{2}(\Omega^{\prime})}\leq C\|u^{\prime}\|_{H^{1}(\Omega)}$ (41)
where we’ve used the fact that for harmonic functions, interior regularity implies that $u^{\prime}\in H^{2}(\Omega^{\prime})$ given that $u^{\prime}\in H^{1}(\Omega)$.
Next, we estimate $\|u^{\prime}\|_{H^{1}(\Omega)}$. By the variational theory of elliptic equations and Lemma 4.2, considering that $u^{\prime}$ satisfies equation (35) with Dirichlet condition on the internal boundary $\Gamma$ and Neumann condition on the external boundary $\Sigma$, we have
$\|u^{\prime}\|_{H^{1}(\Omega)}\leq C\|u^{\prime}\|_{H^{1/2}(\Gamma)}$ (42)
@@
Noting the boundary condition $u^{\prime}|_{\Gamma}=V_{n}\partial_{n}u$ on the internal boundary $\Gamma$, we have
$\|u^{\prime}\|_{H^{1/2}(\Gamma)}=\|V_{n}\partial_{n}u\|_{H^{1/2}(\Gamma)}$ (43)
@@
$\|V_{n}\partial_{n}u\|_{H^{1/2}(\Gamma)}\leq C\|V_{n}\|_{C^{1}(\Gamma)}\|\partial_{n}u\|_{H^{1/2}(\Gamma)}$ (44)
Since under the regularity assumption (39), $\partial_{n}u\in H^{7/2}(\Gamma)\hookrightarrow H^{1/2}(\Gamma)$ and the norm is bounded, i.e., there exists a constant $C>0$, independent of $V(0)$, such that $\|\partial_{n}u\|_{H^{1/2}(\Gamma)}\leq C$, we obtain
$\|V_{n}\partial_{n}u\|_{H^{1/2}(\Gamma)}\leq C\|V_{n}\|_{C^{1}(\Gamma)}$ (45)
Combining the above inequalities, we have
$\max_{x\in\Sigma}|u^{\prime}(x)|\leq C\|u^{\prime}\|_{C^{0}(\Sigma)}\leq C\|u^{\prime}\|_{H^{1}(\Omega)}\leq C\|u^{\prime}|_{\Gamma}\|_{H^{1/2}(\Gamma)}\leq C\|V_{n}\|_{C^{1}(\Gamma)}$ (46)
@@
$|dF(\Gamma;V(0))|\leq\max_{x\in\Sigma}|u^{\prime}(x)|\leq C\|V_{n}\|_{C^{1}(\Gamma)}\leq C\|V(0)\|_{C^{1}(\Gamma;\mathbb{R}^{3})}$ (47)
@@
Next, from (33)
$|d^{2}F(\Gamma;V_{1}(0);V_{2}(0))|=$ $|u^{\prime\prime}(x_{0})|\leq\max_{x_{\in}\Sigma}|u^{\prime\prime}(x)|\leq C\|u^{\prime\prime}\|_{H^{1/2}(\Gamma)}$ (48)
$=$ $C\|\psi\|_{H^{1/2}(\Gamma)}$ (49)
$=$ $\|(V_{1,n}V_{2,n}H-V_{1,\tau}\cdot(DnV_{2,\tau}))\partial_{n}u\|_{H^{1/2}(\Gamma)}$ (50)
$+$ $\|V_{1,n}\partial_{n}u_{2}^{\prime}\|_{H^{1/2}(\Gamma)}+\|V_{2,n}\partial_{n}u_{1}^{\prime}\|_{H^{1/2}(\Gamma)}$ (51)
$+$ $\|(V_{1,\tau}(\tau\cdot\nabla V_{2,n})\partial_{n}u\|_{H^{1/2}(\Gamma)}+\|V_{2,\tau}(\tau\cdot\nabla V_{1,n})\partial_{n}u\|_{H^{1/2}(\Gamma)}$ (52)
$=$ $I_{1}+I_{2}+I_{3}+I_{4}+I_{5}.$ (53)
@@
$\leq$ $C\|(|V_{1,n}V_{2,n}|+|V_{1,\tau}V_{2,\tau}|)\partial_{n}u\|_{H^{1/2}(\Gamma)}.$ (55)
@@
Applying this to our case, and noting that $\partial_{n}u\in H^{7/2}(\Gamma)\subset H^{1/2}(\Gamma)$ with bounded norm due to (39), we get:
$\|(|V_{1,n}V_{2,n}|+|V_{1,\tau}V_{2,\tau}|)\partial_{n}u\|_{H^{1/2}(\Gamma)}$ $\leq C\||V_{1,n}V_{2,n}|+|V_{1,\tau}V_{2,\tau}|\|_{C^{1}(\Gamma)}\|\partial_{n}u\|_{H^{1/2}(\Gamma)}$ (57)
$\leq C\||V_{1,n}V_{2,n}|+|V_{1,\tau}V_{2,\tau}|\|_{C^{1}(\Gamma)}$ (58)
@@
$\||V_{1,n}V_{2,n}|+|V_{1,\tau}V_{2,\tau}|\|_{C^{1}(\Gamma)}$ $\leq\|V_{1,n}V_{2,n}\|_{C^{1}(\Gamma)}+\|V_{1,\tau}V_{2,\tau}\|_{C^{1}(\Gamma)}$ (59)
@@
The components $V_{j,n}$ and $V_{j,\tau}$ are projections of $V_{j}$ onto the normal and tangential directions, respectively. These projections involve the geometry of the boundary through the normal vector $n$ and tangent vectors. Since the boundary $\Gamma$ is $C^{5}$, the normal and tangent vectors are at least $C^{4}$. When we project a vector field onto these directions and take derivatives, we need to account for derivatives of both the vector field and the direction vectors.
For a vector field $V_{j}\in C^{k}(\Gamma;\mathbb{R}^{3})$, its normal component $V_{j,n}=V_{j}\cdot n$ satisfies:
$\|V_{j,n}\|_{C^{1}(\Gamma)}\leq C\|V_{j}\|_{C^{2}(\Gamma;\mathbb{R}^{3})}$ (61)
This is because computing $\nabla_{\Gamma}(V_{j}\cdot n)$ involves both $\nabla_{\Gamma}V_{j}$ and $\nabla_{\Gamma}n$, requiring one more degree of smoothness.
@@
$\|V_{j,\tau}\|_{C^{1}(\Gamma)}\leq C\|V_{j}\|_{C^{2}(\Gamma;\mathbb{R}^{3})}$ (62)
Therefore:
$\||V_{1,n}V_{2,n}|+|V_{1,\tau}V_{2,\tau}|\|_{C^{1}(\Gamma)}$ $\leq C(\|V_{1}\|_{C^{2}(\Gamma;\mathbb{R}^{3})}\|V_{2}\|_{C^{2}(\Gamma;\mathbb{R}^{3})})$ (63)
$\leq C\|V_{1}\|_{C^{2}(\Gamma;\mathbb{R}^{3})}\|V_{2}\|_{C^{2}(\Gamma;\mathbb{R}^{3})}$ (64)
@@
The terms $I_{2}$ through $I_{5}$ involve higher derivatives of the vector fields and the boundary, and their analysis requires even higher regularity. In particular, terms like $V_{j,\tau}\cdot\nabla V_{k,n}$ involve third derivatives of the vector fields, which is why we ultimately need:
$I_{j}\leq C\|V_{1}\|_{C^{3}(\Gamma;\mathbb{R}^{3})}\|V_{2}\|_{C^{3}(\Gamma;\mathbb{R}^{3})},\quad j=1,2,3,4,5$ (66)

### v2

@@ 3.2 Generative Prior-Guided Optimization on Learned Shape Manifolds
Following [42], we assume the regularity condition that $\nabla_{x}f_{\theta}(z,x)\neq 0$ whenever $f_{\theta}(z,x)=0$, ensuring well-defined surfaces with consistent normal vectors. This condition is naturally satisfied by properly trained neural SDF models that accurately approximate true signed distance functions.
The pre-trained 3D generative model, trained on representative shape datasets, induces a low-dimensional manifold capturing the distribution of plausible geometries. Optimization in the latent space constitutes a principled search over this manifold to best explain the measurements. This provides exceptionally potent regularization that constrains solutions to be anatomically plausible, effectively addressing the ill-posedness inherent in EIT reconstruction.
This latent representation fundamentally transforms the optimization landscape by dramatically reducing problem dimensionality — from the infinite-dimensional space of all possible interfaces to a compact latent space, typically of dimension $d\approx 256\ll\infty$. Such dimensionality reduction not only enhances computational efficiency but also provides implicit regularization that helps mitigate the severe ill-posedness inherent in EIT reconstruction. The resulting optimization process can be interpreted as a projection of physics-based gradients onto the learned shape manifold, where physics dictates the descent direction while the generative prior constrains the path to anatomically plausible solutions.
Figure 3: Integration of a pre-trained 3D generative prior with EIT reconstruction: (a) The generative model encodes plausible shapes into a low-dimensional manifold $\mathcal{M}$. (b) Our physics-informed optimization in $\mathcal{M}$ minimizes the EIT data misfit while constraining solutions to the learned manifold.
3.3 Adjoint-Based Gradients in Latent Space
To solve the optimization problem (13) efficiently, we derive analytical expressions for the gradient of the objective function with respect to the latent variables. This requires extending classical shape calculus to optimization on a generative latent manifold. Let $V:\mathbb{R}\times\mathbb{R}^{3}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\rightarrow}\mathbb{R}^{3}$ be a velocity field and let $x(t,X)$ be the solution to
@@ 3.3 Adjoint-Based Gradients in Latent Space
Let $\Omega_{z}$ be the domain bounded externally by $\Sigma$ and internally by $\Gamma_{z}$, and assume that $\Omega_{z}$ is a bounded domain of class $C^{2}$. Let $\mathcal{L}(z)$ be defined in (13). If $\nabla_{x}f_{\theta}$ does not vanish on $\Gamma_{z}$, and $n=-\frac{\nabla_{x}f_{\theta}}{\|\nabla_{x}f_{\theta}\|}$ denotes the outward normal to $\Omega_{z}$ on $\Gamma_{z}$, then:
@@
This gradient formulation provides a rare mathematical bridge between the physics-driven world of shape derivatives and the data-driven, probabilistic world of generative models. The integrand in (16) elegantly combines three critical components: (1) the sensitivity of the forward solution $\frac{\partial u}{\partial n}$, (2) the sensitivity of the adjoint solution $\frac{\partial w}{\partial n}$, and (3) the shape-to-latent mapping $\frac{\nabla_{z}f_{\theta}}{\|\nabla_{x}f_{\theta}\|}$, linking variations in data-misfit to perturbations in the learned shape manifold. This formulation provides a rigorous bridge between physics-driven shape derivatives and data-driven generative latent variables, enabling efficient navigation of the learned manifold toward solutions that match the measurements.
Building on this gradient expression, we can update $z$ using stochastic gradient descent algorithms. In our experiments, we adopt the Adam optimizer, which efficiently combines gradient and momentum information, although other gradient-based methods could also be employed. The complete algorithm description is presented in Algorithm 1. Our approach offers several key advantages, including efficient low-dimensional optimization enabled by the pre-trained generative prior. This combination enhances overall performance, particularly for challenging 3D high-contrast EIT reconstruction problems where maintaining correct topology is crucial for accurate interpretation.
Algorithm 1 Generative Prior-Guided Neural Interface Reconstruction for 3D EIT
1:Input: Voltage measurements, pre-trained 3D generative model $f_{\theta}$.
2:Initialize: Latent code $z_{0}$, learning rates $\{\alpha_{k}\}$, max iterations NN.
3:for $k=0$ to NN do
@@ 4 Convergence Analysis
$d^{2}J(\Gamma;V_{1};V_{2})=\lim_{t\rightarrow 0^{+}}\frac{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}dJ(\Gamma_{t}(V_{2});V_{1}(t))}-dJ(\Gamma;V_{1}(0))}{t}.$ (18)
Following [10, Theorem 6.2, Chapter 9], we have
@@
$d^{2}J(\Gamma;V_{1}(0);V_{2}(0))$ $:=\lim_{t\rightarrow 0^{+}}\frac{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}dJ(T_{t}(\Gamma,V_{2});V_{1}(0))-dJ(\Gamma;V_{1}(0))}}{t},$ (20)
@@
For a detailed discussion of second-order shape derivatives, we refer to [10, 39] and [21, Appendix].
@@
Let $\Gamma$ be of class $C^{3}$. Then there exists $C>0$ independent of $V$ such that:
$|dJ(\Gamma;V(0))|$ $\leq C\|V(0)\|_{C^{1}(\Gamma;\mathbb{R}^{3})},\quad\forall V\in C([0,\varepsilon);C^{1}(\overline{\Omega},\mathbb{R}^{3}))$ (24)
$|d^{2}J(\Gamma;V_{1}(0);V_{2}(0))|$ $\leq C\|V_{1}(0)\|_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}C^{2}}(\Gamma,\mathbb{R}^{3})}\|V_{2}(0)\|_{C^{2}(\Gamma,\mathbb{R}^{3})},$ (25)
for all $V_{1},V_{2}\in C([0,\varepsilon);C^{2}(\overline{\Omega},\mathbb{R}^{3}))$, where $\varepsilon>0$ is sufficiently small.
@@
Let $\Omega\subset\mathbb{R}^{n}$ be a bounded Lipschitz domain representing the conducting region (i.e., $\Omega=D\setminus\overline{S}$), with its boundary explicitly partitioned into two disjoint parts: $\partial\Omega=\Gamma\cup\Sigma$. We consider the following auxiliary boundary value problem restricted strictly to the domain $\Omega$:
@@
By integration by parts, we can easily get
$\int_{\Omega}|\nabla u|^{2}dx=\int_{\Gamma}f\partial_{n}u\,dS.$ (28)
Then using Cauchy-Schwarz inequality and trace theorem, we have
@@
The generalized Poincaré inequality incorporating boundary trace norms[16] gives that
@@
which maps an admissible boundary $\Gamma$ to the value of the state function $u$ at $x_{0}\in\Sigma$. By [1, Theorem 2], $F$ is twice continuously differentiable with
@@
with the boundary data
@@
$+(V_{1,\tau}\cdot\nabla V_{2,n}+V_{2,\tau}\cdot\nabla V_{1,n})\partial_{n}u,$ (34)
where $u$ is the solution of the state problem, and $n$ denotes the unit outward normal to the boundary $\Gamma$. The notation $V_{j,n}=V_{j}(0)\cdot n$ and $V_{j,\tau}=V_{j}(0)-V_{j,n}n$ denote the normal scalar component and tangential vector projection of the velocity field $V_{j}(0)$, respectively. $H$ denotes the mean curvature of the boundary $\Gamma$, and $Dn$ represents the derivative of the normal vector field on $\Gamma$. Moreover, $u_{j}^{\prime}$ for $j=1,2$, is the solution to the following boundary value problem
$\left\{\begin{aligned} \Delta u_{j}^{\prime}&=0\quad\text{ in }\Omega,\\ u_{j}^{\prime}&={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}-V_{j,n}\partial_{n}u}\quad\text{ on }\Gamma,\\ \partial_{n}u_{j}^{\prime}&=0\quad\text{ on }\Sigma.\end{aligned}\right$. (35)
Still by [1, Theorem 1], we have
$\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}dF(\Gamma;V_{j}(0))=u_{j}^{\prime}(x_{0}),\quad j=1,2.$} (36)
@@
$|dF(\Gamma;V(0))|$ $\leq C\|V(0)\|_{C^{1}(\Gamma;\mathbb{R}^{3})},\quad\forall V\in C([0,\varepsilon);C^{1}(\overline{\Omega},\mathbb{R}^{3})),$ (37)
$|d^{2}F(\Gamma;V_{1}(0);V_{2}(0))|$ $\leq C\big\|V_{1}(0)\|_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}C^{2}}(\Gamma,\mathbb{R}^{3})}\|V_{2}(0)\|_{C^{2}(\Gamma,\mathbb{R}^{3})},\,\quad\forall V_{1},V_{2}\in C([0,\varepsilon);C^{2}(\overline{\Omega},\mathbb{R}^{3})).$ (38)
@@
First, the $C^{3}$ regularity of the boundary $\Gamma$ implies that, for a bounded neighborhood $\Omega^{+}$ of $\Gamma$ within $\Omega$,
$u\in H^{3}(\Omega^{+}),\quad u|_{\Gamma}\in H^{5/2}(\Gamma)\quad\mbox{and}\quad\partial_{n}u\in H^{3/2}(\Gamma).$ (39)
@@
$|dF(\Gamma;V(0))|=|u^{\prime}(x_{0})|\leq\max_{x\in\Sigma}|u^{\prime}(x)|=\|u^{\prime}\|_{C^{0}(\Sigma)}.$ (40)
@@
$\|u^{\prime}\|_{C^{0}(\Sigma)}\leq C\|u^{\prime}\|_{H^{3/2}(\Sigma)}\leq C\|u^{\prime}\|_{H^{2}(\Omega^{\prime})}\leq C\|u^{\prime}\|_{H^{1}(\Omega)},$ (41)
where we have used local boundary elliptic regularity. Since $u^{\prime}$ is harmonic in $\Omega$ and satisfies the homogeneous Neumann boundary condition $\partial_{n}u^{\prime}=0$ on the smooth boundary $\Sigma$, the regularity is elevated locally near the boundary, guaranteeing that $u^{\prime}\in H^{2}(\Omega^{\prime})$ given $u^{\prime}\in H^{1}(\Omega)$.
Next, we estimate $\|u^{\prime}\|_{H^{1}(\Omega)}$. By the variational theory of elliptic equations, Lemma 4.2, and the fact that $u^{\prime}$ satisfies equation (35) with the boundary condition $u^{\prime}|_{\Gamma}=-V_{n}\partial_{n}u$, we obtain
$\|u^{\prime}\|_{H^{1}(\Omega)}\leq C\|u^{\prime}\|_{H^{1/2}(\Gamma)}=C\|-V_{n}\partial_{n}u\|_{H^{1/2}(\Gamma)}=C\|V_{n}\partial_{n}u\|_{H^{1/2}(\Gamma)},$ (42)
@@
$\|V_{n}\partial_{n}u\|_{H^{1/2}(\Gamma)}\leq C\|V_{n}\|_{C^{1}(\Gamma)}\|\partial_{n}u\|_{H^{1/2}(\Gamma)}.$ (43)
Since under the regularity assumption (39), $\partial_{n}u\in{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}H^{3/2}}(\Gamma)\hookrightarrow H^{1/2}(\Gamma)$ and the norm is bounded, i.e., there exists a constant $C>0$, independent of $V(0)$, such that $\|\partial_{n}u\|_{H^{1/2}(\Gamma)}\leq C$, we obtain
$\|V_{n}\partial_{n}u\|_{H^{1/2}(\Gamma)}\leq C\|V_{n}\|_{C^{1}(\Gamma)}.$ (44)
Combining the above inequalities, especially the local boundary regularity(41), we have:
$\max_{x\in\Sigma}|u^{\prime}(x)|\leq C\|u^{\prime}\|_{C^{0}(\Sigma)}\leq C\|u^{\prime}\|_{H^{1}(\Omega)}\leq C\|u^{\prime}|_{\Gamma}\|_{H^{1/2}(\Gamma)}\leq C\|V_{n}\|_{C^{1}(\Gamma)},$ (45)
@@
$|dF(\Gamma;V(0))|\leq\max_{x\in\Sigma}|u^{\prime}(x)|\leq C\|V_{n}\|_{C^{1}(\Gamma)}\leq C\|V(0)\|_{C^{1}(\Gamma;\mathbb{R}^{3})}.$ (46)
@@
Next, from (33), by applying the exact same local boundary regularity and embedding arguments (cf. (41)) as we did for $u^{\prime}$, we establish the continuous chain of bounds for $u^{\prime\prime}$:
$|d^{2}F(\Gamma;V_{1}(0);V_{2}(0))|=$ $|u^{\prime\prime}(x_{0})|\leq\|u^{\prime\prime}\|_{C^{0}(\Sigma)}$ (47)
$\leq$ $C\|u^{\prime\prime}\|_{H^{1/2}(\Gamma)}=C\|\psi\|_{H^{1/2}(\Gamma)}$
$\leq$ $C\big(\|(V_{1,n}V_{2,n}H-V_{1,\tau}\cdot(DnV_{2,\tau}))\partial_{n}u\|_{H^{1/2}(\Gamma)}$
$\quad+\|V_{1,n}\partial_{n}u_{2}^{\prime}\|_{H^{1/2}(\Gamma)}+\|V_{2,n}\partial_{n}u_{1}^{\prime}\|_{H^{1/2}(\Gamma)}$
$\quad+\|(V_{1,\tau}\cdot\nabla V_{2,n})\partial_{n}u\|_{H^{1/2}(\Gamma)}+\|(V_{2,\tau}\cdot\nabla V_{1,n})\partial_{n}u\|_{H^{1/2}(\Gamma)}\big)$
$=$ $C(I_{1}+I_{2}+I_{3}+I_{4}+I_{5}).$
@@
$\leq$ $C\|(|V_{1,n}V_{2,n}|+|V_{1,\tau}||V_{2,\tau}|)\partial_{n}u\|_{H^{1/2}(\Gamma)},$ (49)
where we used the fact that the shape operator $Dn$ is bounded on the smooth boundary $\Gamma$.
@@
Applying this to our case, and noting that $\partial_{n}u\in H^{3/2}(\Gamma)\hookrightarrow H^{1/2}(\Gamma)$ with bounded norm due to (39), we get:
$\|(|V_{1,n}V_{2,n}|+|V_{1,\tau}||V_{2,\tau}|)\partial_{n}u\|_{H^{1/2}(\Gamma)}$ $\leq C\||V_{1,n}V_{2,n}|+|V_{1,\tau}||V_{2,\tau}|\|_{C^{1}(\Gamma)}\|\partial_{n}u\|_{H^{1/2}(\Gamma)}$ (51)
$\leq C\||V_{1,n}V_{2,n}|+|V_{1,\tau}||V_{2,\tau}|\|_{C^{1}(\Gamma)}$ (52)
@@
$\||V_{1,n}V_{2,n}|+|V_{1,\tau}||V_{2,\tau}|\|_{C^{1}(\Gamma)}$ $\leq\|V_{1,n}V_{2,n}\|_{C^{1}(\Gamma)}+\||V_{1,\tau}||V_{2,\tau}|\|_{C^{1}(\Gamma)}$ (53)
@@
The components $V_{j,n}$ and $V_{j,\tau}$ are projections of $V_{j}$ onto the normal and tangential directions, respectively. Specifically, the tangential projection is given by $V_{j,\tau}=V_{j}-V_{j,n}n$. This projection relies entirely on the geometry of the boundary through the normal vector $n$. Since the boundary $\Gamma$ is $C^{3}$, the normal vector field is at least $C^{2}$. When we take derivatives of this projection, we only need to account for the derivatives of the vector field and the normal vector $n$.
For a vector field $V_{j}\in C^{1}(\Gamma;\mathbb{R}^{3})$, its normal component $V_{j,n}=V_{j}\cdot n$ satisfies:
$\|V_{j,n}\|_{C^{1}(\Gamma)}\leq C\|V_{j}\|_{C^{1}(\Gamma;\mathbb{R}^{3})}.$ (55)
This holds because computing the surface gradient $\nabla_{\Gamma}(V_{j}\cdot n)=(\nabla_{\Gamma}V_{j})n+(\nabla_{\Gamma}n)V_{j}$ involves only the first-order derivatives of the vector field $V_{j}$ and the first-order derivatives of the normal vector $n$ (which requires the boundary $\Gamma$ to be at least $C^{2}$).
@@
$\|V_{j,\tau}\|_{C^{1}(\Gamma)}\leq C\|V_{j}\|_{C^{1}(\Gamma;\mathbb{R}^{3})}.$ (56)
Therefore, utilizing the fact that $\|\cdot\|_{C^{1}}\leq\|\cdot\|_{C^{2}}$, we can bound the product as:
$\||V_{1,n}V_{2,n}|+|V_{1,\tau}||V_{2,\tau}|\|_{C^{1}(\Gamma)}$ $\leq C\|V_{1}\|_{C^{1}(\Gamma;\mathbb{R}^{3})}\|V_{2}\|_{C^{1}(\Gamma;\mathbb{R}^{3})}$
$\leq C\|V_{1}\|_{C^{2}(\Gamma;\mathbb{R}^{3})}\|V_{2}\|_{C^{2}(\Gamma;\mathbb{R}^{3})}.$ (57)
@@
The terms $I_{2}$ through $I_{5}$ involve at most first-order derivatives of the vector fields and the normal vector. Bounding their $H^{1/2}(\Gamma)$ norms via the product rule requires one additional derivative. Consequently, $C^{2}$ regularity for the vector fields and $C^{3}$ regularity for the boundary $\Gamma$ are sufficient to yield:
$I_{j}\leq C\|V_{1}\|_{C^{2}(\Gamma;\mathbb{R}^{3})}\|V_{2}\|_{C^{2}(\Gamma;\mathbb{R}^{3})},\quad j=2,3,4,5.$ (59)